OpenStudy (anonymous):

Please help for a medal and fan. 1. A square has an area of x² - 8x + 16. A rectangle has a length that is two times the side of the square. The width of the rectangle is 3 centimeters less than the side of the square. What are the dimensions of the rectangle? Length = ____ centimeters Width = ____ centimeters 2. Christian is framing an Algebra certificate with a frame that has a 1-inch border. If the area of the frame is 396 square inches, and the length of the certificate is 4 inches greater than its width, what are the dimensions of the certificate, excluding the frame?

OpenStudy (surana):

@Hero ?

OpenStudy (surana):

This isn't something I'm really good with, sorry.

OpenStudy (anonymous):

do u go to connections 8 grade @AMYCARTER

OpenStudy (nikato):

do you know how to factor x² - 8x + 16

OpenStudy (anonymous):

yes @nikato it's \[(x-4)^{2}\]

OpenStudy (nikato):

yes. and what is x?

OpenStudy (anonymous):

I'm not sure...

OpenStudy (nikato):

you want to set whats in the parentheses equal to 0 so x-4=0 so x=?

OpenStudy (anonymous):

do u kgo to conneciton 8grade

OpenStudy (anonymous):

x would equal 4 then

OpenStudy (nikato):

correct. so that would be the length of a side of a square. remember all 4 sides of a square are equal so they all =4

OpenStudy (nikato):

let s= side of sqaure= 4 rectangle has a length that is two times the side of the square. so length= 2s =2(4) =?

OpenStudy (anonymous):

2 times 4 is 8

OpenStudy (nikato):

yes and width of the rectangle is 3 centimeters less than the side of the square width= s-3 =4-3 =?

OpenStudy (anonymous):

1 @nikato

OpenStudy (anonymous):

okay so for the first one 8,1. How about the second one @nikato?

OpenStudy (nikato):

one moment please as i think about the problem

OpenStudy (anonymous):

okay thank you so much for coming to help me @nikato

OpenStudy (nikato):